OpenStudy (anonymous):

A tennis ball on Mars, where the acclertation due to gravity is 0.379g, and air resistance is negligible, is hit directly upwards and returns to the same level at 8.5 seconds time. A) how high above its original point did the ball go? B) what was it initial velocity after it was hit?

OpenStudy (anonymous):

MY variables are a= 3.72 ms^2 T=8.5 seconds Vf= 0 I converted acceleration to m/s^2 and got 3.72 m/s^2 and i assumed Vyf is =0

OpenStudy (anonymous):

@sammixboo do you have time to take a look?

OpenStudy (anonymous):

using the kinematic equation d=vit+1/2at^2 i got 33.6 as my distance

OpenStudy (anonymous):

@ganeshie8 am i attacking this problem the right way?

ganeshie8 (ganeshie8):

8.5 seconds is the total time for going up and falling all the way back again so you need to take t = 8.5/2 for finding the maximum height

OpenStudy (anonymous):

to get my distance, started from the top, which gave me my variables vi=0 a=3.72 t=4.25 i use \[\delta d= 67.13\] and \[\delta t = 8.5\] i use the equation d= vit+1/2at^2 but i get a vi of 67.13 ms which doesnt seem right

ganeshie8 (ganeshie8):

you may start by finding the initial velocity

ganeshie8 (ganeshie8):

\(v=u+at \) plugin \(v = 0\), \(a = -3.72\) and \(t = 4.25\)

ganeshie8 (ganeshie8):

\(0 = u - 3.72 (4.25)\) \(u = ?\)

OpenStudy (anonymous):

A) 33.6 b)15.81

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

thanks, it seems correct now!

ganeshie8 (ganeshie8):

yw